A firm buys 3 shipments of parts each month. The purchasing agent selects at random from among four in-state suppliers and six out-of-state suppliers. What is the probability that orders are placed with
a) The in-state suppliers only
b) The out-of-state suppliers only
c) At least one in- state suppliers
1
Expert's answer
2020-08-13T17:07:58-0400


in-state suppliers (I) "=4"

Out of state suppliers (O) "=6"

Suppliers picked (r) "=3"

Total suppliers (N) "=10"


Probability of selecting 3 suppliers from the total



"=\frac{_I C_x * _OC_y}{_N C_r}"

Where 'I' are the in-state suppliers selected and 'O' the number of out of state suppliers


solution a) in-state suppliers only



"=\frac{4 C_3}{10C_3} = \frac{4}{120}"

answer: 0.0333


solution b) out of state suppliers only

"=\frac{6C_3}{10C_3} = \frac{20}{120}"

answer: 0.1667


solution c) atleast 1 in-state supplier


when in-state supplier "=1"



"Probability =\frac{4C_1 * 6C_2}{10C_3} = \frac{4*15}{120} = 0.5"

when in-state suppliers "=2"



"Probability = \frac{4C_2 * 6C_1}{10C_3} = \frac{6*6}{120}=0.3"

when in-state suppliers"=3"


"Probability = 0.0333"


Probability of atleast 1 in-state supplier



"=0.5+0.3+0.0333"

answer: 0.8333